# VMarker-Pro: Probabilistic 3D Human Mesh Estimation from Virtual Markers

### v3

@@ 4.1 Datasets and metrics
H3.6M [13]. We use (S1, S5, S6, S7, S8) for training and (S9, S11) for testing. For the deterministic estimation task, as in [6, 1, 87, 89], we report Mean Per Joint Position Error (MPJPE) and PA-MPJPE for poses that are derived from the estimated meshes. We also report Mean Per Vertex Error (MPVE) for the whole mesh. For the probabilistic estimation task, we report the minimum errors over $S$ hypotheses to measure the estimated distribution accuracy following a standard practice in recent works [43, 45].
3DPW [14] is collected in natural scenes. Following previous works [87, 89, 47, 40, 43, 45], we use the train set of 3DPW to learn the model and evaluate on the test set. The same evaluation metrics as H3.6M are used. We further evaluate the performance of our probabilistic framework VMarker-Pro on two subsets of 3DPW, i.e.  3DPW-OC [14, 95] and 3DPW-PC [14, 88] which are composed by object- and person-specific occlusion, respectively. To ensure a fair comparison, as suggested by [96], we do not use the 3DPW training set in these particular evaluations.
SURREAL [27] is a large-scale synthetic dataset with GT SMPL annotations and has diverse samples in terms of body shapes, backgrounds, etc.  We use its training set to train a model and evaluate the test split following [1]. The same evaluation metrics as H3.6M [13] are reported.
@@ 4.2 Implementation Details
We learn $64$ virtual markers on the H3.6M [13] training set. We use the same set of markers for all datasets instead of learning a separate set for each one. Following [6, 1, 17, 40, 16, 84, 89, 87], we conduct mix-training by using MPI-INF-3DHP [97], UP-3D [15], and COCO [98] training set for experiments on the H3.6M and 3DPW datasets.
We adapt a 3D pose estimator [18] with HRNet-W48 [99] as the image feature backbone for estimating the 3D virtual markers. We set the number of voxels in each dimension to be $64$, i.e.  $D=H=W=64$ for 3D heatmaps. Following [6, 16, 17], we crop every single human region from the input image and resize it to $256\times 256$. The dimensions of the 2D feature map are $64$, i.e.  $H_{b}=W_{b}=64$. The 2D feature map has $C_{b}=48$ feature channels. The processed $\mathbf{F}^{2d}$ feature channels are set to be $C=64$. The denoiser network has $B=3$ blocks. To train VMarker, we use Adam [100] optimizer to train the whole framework for $40$ epochs with a batch size of $32$. The learning rates for the 3D VM estimation branch and the updating branch are set to $5\times 10^{-4}$ and $1\times 10^{-3}$, respectively, which are decreased by half after the $30^{th}$ epoch. To train VMarker-Pro, we use Adam [100] optimizer to train the whole framework for $50$ epochs with a batch size of $80$. The learning rates for the 3D VM estimation branch, the updating branch, and the denoiser are set to $4\times 10^{-5}$, $2\times 10^{-4}$, and $2\times 10^{-3}$, respectively. They are decayed by $0.5$, $0.5$, and $0.1$ after $30^{th}$ and $40^{th}$ epochs, respectively. For inference, we employ DDIM [60] strategy and generate each hypothesis in $T=10$ steps with $\eta=0$. Please check the supplementary for more details.
TABLE IV: Comparison to the state-of-the-arts on SURREAL [27] dataset. means training on the test split with 2D supervisions. “Skel. + Seg.” means using skeleton and segmentation together. $S$ is the number of hypotheses.
@@
HMR [6] - - 85.1 73.6 55.4
BodyNet [46] Skel. + Seg. - 65.8 - -
GraphCMR [16] 3D vertices - 103.2 87.4 63.2
SPIN [33] - - 82.3 66.7 43.7
DecoMR [38] IUV image - 68.9 52.0 43.0
Pose2Mesh [1] 3D skeleton - 68.8 56.6 39.6
PC-HMR [36] 3D skeleton - 59.8 51.7 37.9
DynaBOA [101] - - 70.7 55.2 34.0
VMarker (Ours) Virtual marker - 44.7 36.9 28.9
10 44.3 36.6 28.0
VMarker-Pro 25 42.0 34.8 26.2
(Ours) 100 38.1 31.0 23.5
Virtual marker 200 36.5 29.8 22.4
TABLE V: Comparison to the state-of-the-art methods on the challenging occlusion-specific 3DPW-OC [14, 95] and 3DPW-PC [14, 88] datasets. The top and bottom blocks are deterministic and probabilistic methods, respectively.
Method Venue Hypothesis 3DPW-OC [14, 95] 3DPW-PC [14, 88]
[…]
SPIN [33] ICCV’19 - 121.4 95.5 60.7 159.8 122.1 77.4
I2L-MeshNet [17] ECCV’20 - 129.5 92.0 61.4 160.2 117.3 80.0
PyMAF [37] ICCV’21 - 113.7 89.6 59.1 154.6 117.5 74.5
ROMP [88] ICCV’21 - - 91.0 62.0 152.8 117.9 79.7
PARE [47] ICCV’21 - 101.5 83.5 57.0 122.4 96.8 64.5
OCHMR [90] CVPR’22 - 145.9 112.2 75.2 149.6 117.5 77.1
3DCrowdNet [91] CVPR’22 - 101.5 83.5 57.1 114.8 90.9 64.4
JOTR [96] ICCV’23 - 92.6 75.7 52.2 109.7 86.5 58.3
VMarker (Ours) [28] CVPR’23 - 92.8 78.4 49.0 110.8 93.9 66.0
HMDiff [45] ICCV’23 25 - - - 143.1 114.2 73.5
10 90.4 77.3 48.6 109.4 89.8 60.2
25 85.4 73.3 46.0 102.9 84.7 56.6
100 77.7 67.1 42.5 94.0 78.0 52.0
VMarker-Pro (Ours) - 200 75.2 64.9 40.9 90.3 74.6 49.7
@@ 4.3.1 Deterministic Track
Results on H3.6M [13]. Table II compares our approach to the deterministic state-of-the-art (SOTA) methods on the H3.6M dataset. Our method achieves superior performance. In particular, it outperforms the methods that use skeletons (Pose2Mesh [1], DSD-SATN [48]), body markers (THUNDR) [40], or IUV image [38, 37] as proxy representations, demonstrating the effectiveness of the virtual marker representation.
@@
Results on 3DPW [14]. We compare our method to the deterministic SOTA methods on the 3DPW dataset in Table II. Our approach achieves SOTA results among all the methods, validating the advantages of the virtual marker representation over the skeleton representation used in Pose2Mesh [1], DSD-SATN [48], and other representations like IUV image used in PyMAF [37]. In particular, our approach outperforms I2L-MeshNet [17], METRO [87], and Mesh Graphormer [89] by a notable margin, suggesting that virtual markers are more suitable and effective representations than using all vertices as most of them are not discriminative enough to be accurately detected.
@@
Results on SURREAL [27]. This dataset has more diverse samples in terms of body shapes. The results are shown in Table IV. Our approach VMarker, outperforms the SOTA methods by a notable margin, especially in terms of MPVE. Figure 1 (top) shows some challenging cases without cherry-picking. The skeleton representation loses the body shape information so the method [1] can only recover mean shapes. In contrast, our approach generates much more accurate mesh estimation results.
4.3.2 Probabilistic Track
Results on H3.6M [13] and 3DPW [14]. Table III presents a comparison between VMarker-Pro and other SOTA probabilistic methods on the H3.6M [13] and 3DPW [14] datasets. Our method significantly surpasses the existing SOTA approaches in terms of accuracy when estimating data distributions. This serves as a validation of the effectiveness of our probabilistic framework, which benefits from the generation capability of diffusion models. Furthermore, as the number of hypotheses ($S$) increases, our method consistently improves, highlighting the diverse and accurate nature of our approach when modeling data distributions. Notably, compared to HMDiff [45], which is also a diffusion model-based method, our incorporation of 2D image cues and the virtual marker representations contributes to the observed notable improvement.
Results on 3DPW-OC [14, 95] and 3DPW-PC [14, 88]. Table V shows the comparison results of our method against existing methods on the 3DPW-OC and 3DPW-PC datasets. Both 3DPW-OC [14, 95] and 3DPW-PC [14, 88] are subsets of the 3DPW [14] dataset, containing scenarios with occlusions by objects and people, respectively. It can be seen that our deterministic estimation results are comparable to the SOTA performance. When adopting the proposed probabilistic modeling approach, the accuracy of our method in estimating the data distribution is significantly improved, thanks to the enhanced accuracy in 3D virtual marker estimation. Figure 5 presents a comparative visualization between VMarker and VMarker-Pro. It can be seen that the deterministic approach, VMarker, may encounter difficulties under occlusions, e.g.  the wrongly estimated right arm and the incorrect arm length. Conversely, the probabilistic method VMarker-Pro, demonstrates a capability to accurately model realistic human poses, offering feasible estimations for the occluded parts.
Figure 5: Qualitative comparison of VMarker and VMarker-Pro on 3DPW-OC [14, 88] testset. For each mesh, we show a projecting view and a side view (with shadow). VMarker wrongly estimates the right arm when heavy occlusion happens while VMarker-Pro provides two reasonable estimates.
Results on SURREAL [27]. Table IV displays the results of our probabilistic approach on the SURREAL [27] dataset. It can be seen that when $S=10$, our probabilistic framework already surpasses existing approaches. This indicates that our method can accurately model the more diverse data distribution. As the number of hypotheses ($S$) increases, the performance of our model continues to improve.
4.4 Ablation Study on VMarker
Virtual marker representation. We compare our method to two baselines in Table VI. First, in baseline (a), we replace the virtual markers of our method with the skeleton representation. The rest are kept the same as ours (c). Our method achieves a much lower MPVE than the baseline (a), demonstrating that the virtual markers help to estimate body shapes more accurately than the skeletons. In baseline (b), we randomly sample $64$ from the $6890$ mesh vertices as virtual markers. We repeat the experiment five times and report the average number. We can see that the result is worse than ours, which is because the randomly selected vertices may not be expressive enough to reconstruct the other vertices or can not be accurately detected from images as they lack distinguishable visual patterns. The results validate the effectiveness of our learning strategy.
TABLE VI: Ablation study of the virtual marker representation for our approach VMarker on H3.6M [13] and SURREAL [27] datasets. “Skeleton” means the sparse landmark joint representation is used. “Rand virtual marker” means the virtual markers are randomly selected from all the vertices without learning. (c) is our method, where the learned virtual markers are used. MPVE error is reported.
@@ 4.4 Ablation Study on VMarker
Figure 6: Mesh estimation results of different methods on H3.6M [13] test set. Our method VMarker with virtual marker representation gets better shape estimation results than Pose2Mesh [1] which uses skeleton representation. Note the waistline of the body and the thickness of the arm.
@@
Figure 1 (top) shows some qualitative results on the SURREAL test set. The meshes estimated by the baseline which uses skeleton representation, i.e.  Pose2Mesh [1], have inaccurate body shapes. This is reasonable because the skeleton is oversimplified and has very limited capability to recover shapes. Instead, it implicitly learns a mean shape for the whole training dataset. In contrast, the mesh estimated by using virtual markers has much better quality due to its strong representation power and therefore can handle different body shapes elegantly. Figure 6 also shows some qualitative results on the H3.6M test set. For clarity, we also draw the intermediate representation (blue balls) in it.
@@
Number of virtual markers. We evaluate how the number of virtual markers affects estimation quality on H3.6M [13] dataset. Figure 2 visualizes the learned 64 markers. Figure 7 visualizes the learned virtual markers when $K=16,32,96$, which are all located on the body surface and close to the extreme points of the mesh. This is expected as mentioned in Section 3.1. Table VII (GT) shows the mesh reconstruction results when we have GT 3D positions of the virtual markers in objective (Eq. 1). When we increase the number of virtual markers, both mesh reconstruction error, i.e.  MPVE, and the regressed landmark joint error, i.e.  MPJPE, steadily decrease. This is expected because using more virtual markers improves the representation power. However, using more virtual markers cannot guarantee smaller estimation errors when we need to estimate the virtual marker positions from images as in our method. This is because the additional virtual markers may have large estimation errors which affect the mesh estimation result. The results are shown in Table VII (Est). Increasing the number of virtual markers $K$ steadily reduces the MPVE errors when $K$ is smaller than $96$. However, if we keep increasing $K$, the error begins to increase. This is mainly because some of the newly introduced virtual markers are difficult to detect from images and therefore bring errors to mesh estimation.
TABLE VII: Ablation study of the different number of virtual markers ($K$) on H3.6M [13] dataset. (GT) Mesh reconstruction results when GT 3D positions of the virtual markers are used in objective (1). (Est) Mesh estimation results obtained by our method VMarker when we use different numbers of virtual markers ($K$).
$K$ GT Est
@@
Figure 7: Visualization of the learned virtual markers when $K=16,32,96$, from left to right, respectively.
Updating coefficient matrix. We compare our method to a baseline which uses the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$. We show the quality comparison in Figure 8. We can see that the estimated mesh by (a) a fixed coefficient matrix has mostly correct pose and shape but there are some artifacts on the mesh while using the (b) updated coefficient matrix can get better mesh estimation results. As shown in Table VIII, using a fixed coefficient matrix gets larger MPVE and MPJPE errors than using the updated one. This is caused by the estimation errors of virtual markers when occlusion happens, which is inevitable since the virtual markers on the back will be self-occluded by the front body. As a result, inaccurate marker positions would bring large errors to the final mesh estimates if we directly use the fixed matrix.
TABLE VIII: Ablation study of the coefficient matrix for VMarker on H3.6M [13] dataset. “fixed” means using the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ to reconstruct the mesh.
@@
(a) VMarker (fixed) 64.7 51.6
(b) VMarker 58.0 47.3
Figure 8: Mesh estimation comparison results when using (a) fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$, and (b) updated $\mathbf{A}$ in our method VMarker. Please zoom in to better see the details.
TABLE IX: Ablation study of the architecture design for the probabilistic method VMarker-Pro on 3DPW [14] dataset. (a-b) ablates the use of 2D feature $\mathbf{F}$ and 2D VMs $\mathbf{P}^{2d}$ in denoiser network, respectively. Ablation (c) removes the 2D VM $\mathbf{P}^{2d}$ sampling in 2D feature $\mathbf{F}$ and directly transforms the global 2D feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ for the denoiser. In ablation (d), the denoiser directly estimates the noise $\bm{\epsilon}$ following the original formula [26], while keeping the rest design the same. (e) is the proposed VMarker-Pro framework.
[…]
$\mathbf{F}$ $\mathbf{P}^{2d}$ $\mathbf{H}^{2d}$
(a) w/o $\mathbf{F}$ 80.1 68.2
(b) w/o $\mathbf{P}^{2d}$ 80.6 68.8
(c) w/o sampling in $\mathbf{F}$ 80.4 69.2
(d) $\rightarrow\bm{\epsilon}$ 80.6 68.9
(e) VMarker-Pro 76.9 66.1
Figure 9: Mesh estimation comparison results of different ablated methods of VMarker-Pro. (a) We ablate the usage of 2D image features $\mathbf{F}$. (b) We ablate the usage of 2D estimated VM $\mathbf{P}^{2d}$. (c) We remove the 2D VM $\mathbf{P}^{2d}$ sampling in 2D feature $\mathbf{F}$ and directly transform the global 2D feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ for the denoiser. (d) We change the estimation target of the denoiser to the noise $\bm{\epsilon}$ instead of the signal itself. (e) is our proposed VMarker-Pro method. We show the GT mesh and its side view on the far right. For baselines (a-e), we show each hypothesis that is denoised from a zero noise for a fair comparison. Each hypothesis presents a projecting view and a side view.
4.5 Ablation Study on VMarker-Pro
Architecture design of the denoiser. Table IX compares VMarker-Pro to four baselines. We report the metric when predicting $S=10$ hypotheses. In baseline (a) and (b), we remove the 2D feature $\mathbf{F}$ and 2D estimated VMs $\mathbf{P}^{2d}$ input for the denoiser, respectively. It can be seen that the errors increase a lot compared to the full model, which validates the effectiveness of the two guidance. In baseline (c), we remove the local sampling by 2D VMs $\mathbf{P}^{2d}$ on 2D feature $\mathbf{F}$, i.e.  the purple arrow in Figure 4. Instead, we directly transform the global 2D image feature $\mathbf{F}$ to $\mathbf{F}^{2d}$ when feeding into the denoiser and the performance degrades notably. In ablation study (d), we alter the denoiser to estimate the noise $\bm{\epsilon}$ rather than the clean 3D VMs $\mathbf{P}(0)$, adjusting the inference process to align with the original denoising formulation by Ho et al.  [26]. This modification results in a decrease in performance. A similar observation was made in previous research by Tevet et al.  [68], where it is suggested that noise estimation might require substantially larger batch sizes to be effective. Unfortunately, we lack the resources to confirm this hypothesis in our settings.
Figure 9 presents the qualitative comparison of different designs. As observed in the image, the left arm of the subject is nearly fully extended. However, due to depth ambiguity, the estimations from baselines (a-d) align with the 2D perspective, yet inaccuracies arise in the 3D estimation of the left forearm when viewed from the side view. Our method (e), which integrates 2D local features, more accurately captures precise 3D poses, achieving estimations that most closely match the GT. This highlights the significance of our design, which not only utilizes 2D local features $\mathbf{F}$ but also leverages 2D VMs $\mathbf{P}^{2d}$ to locally sample these features effectively.
TABLE X: Results of using different denoising steps $T$ for VMarker-Pro on H3.6M [13] and 3DPW [14] datasets.
[…]
1 386.5 370.4 372.2 358.9 339.6
2 77.0 66.2 104.7 92.6 326.6
4 58.8 48.3 80.0 69.2 301.0
5 57.1 46.9 78.3 67.6 292.6
10 55.8 45.4 77.3 66.6 247.4
15 55.5 45.0 76.4 65.5 208.0
20 55.8 45.3 76.8 65.9 189.0
100 55.6 45.2 76.9 65.9 66.4
Figure 10: Mesh estimation results of using different denoising steps $T$ for VMarker-Pro on a sample data from H3.6M [13] dataset. Each hypothesis is denoised from a zero initial noise for a fair comparison.
Denoising steps. Table X shows the metrics on the H3.6M [13] and 3DPW [14] test sets when using different numbers of denoising steps $T$. We report the metrics of predicting $S=10$ hypotheses. We also report the speed of the whole model inferring a hypothesis on an NVIDIA A100-PCIE-80GB GPU. It can be seen that using only $T=1$ denoising step results in significant errors, and the estimated mesh is not human-shaped, as shown in Figure 10. As the number of denoising steps $T$ increases, the accuracy of the estimated mesh gradually improves, and the error decreases. When $T>=10$, the accuracy of the mesh estimation stabilizes, and further steps yield only marginal improvements, as validated by the marginal reduction in error and the indistinguishable differences in visualization. Therefore, based on the acceleration of DDIM [60], we choose $T=10$ steps for a good balance between estimation performance and inference speed.
TABLE XI: Robustness to occlusion for VMarker and VMarker-Pro on 3DPW [14] test set when different body parts are occluded. We report MPVE error.
Occ. Parts VMarker VMarker-Pro
$S=10$ $S=100$
None 77.9 77.3 67.4
2 Arms 79.2 78.7 68.8
2 Legs 78.3 77.6 67.8
Body 78.6 78.1 68.2
Random 78.7 77.9 68.4
Figure 11: Top: Meshes estimated by our deterministic approach VMarker on images from 3DPW test set. The rightmost case in the dashed box shows a typical failure. Bottom: Meshes estimated by VMarker on Internet images with challenging cases (extreme shapes or in a long dress).
Figure 12: (a) For each virtual marker (represented by a star), we highlight the top 30 most affected vertices (represented by a colored dot) based on average coefficient matrix $\mathbf{A}$. (b) For each vertex (dot), we highlight the top 3 virtual markers (star) that contribute the most. The dependency has a strong locality which improves the robustness when some virtual markers cannot be accurately detected.
Figure 13: Qualitative comparison of VMarker and VMarker-Pro on 3DPW [14] testset. VMarker may fail when facing ambiguity, while VMarker-Pro could generate multiple reasonable solutions. For each mesh estimate, we present both a projected view and a side view in each column. We show three hypotheses generated by VMarker-Pro.
Figure 14: Meshes estimated by VMarker-Pro on Internet images. For each case in a row, we present three hypotheses, showing both a projected view and a side view in each column. The rightmost column shows the overlapped side views to unveil the subtle differences.
Figure 15: A typical failure case of VMarker-Pro. The heavy occlusion and ambiguity cause the model to incorrectly estimate the right arm and right leg of the male dancer.
Robustness to occlusion. We report results of VMarker and VMarker-Pro when different image regions of corresponding virtual markers are occluded by a synthetic random mask in Table XI. The errors of VMarker are slightly larger than the original image (None). We further analyze how inaccurate virtual markers would affect the mesh estimation, i.e.  when part of human body is occluded or truncated. According to the finally learned coefficient matrix $\mathbf{A}$ of our model VMarker, we highlight the relationship weights among virtual markers and all vertices in Figure 12. We can see that our model actually learns local and sparse dependency between each vertex and the virtual markers, e.g.  for each vertex, the virtual markers that contribute the most are in a near range as shown in Figure 12 (b). Therefore, in inference, if a virtual marker has inaccurate position estimation due to occlusion or truncation, the dependent vertices may have inaccurate estimates, while the rest will be barely affected.
This locality enhances the robustness of the VMarker-Pro method against occlusion as well. When sampling $S=10$ hypotheses, VMarker-Pro exhibits a minor increase in error under various circumstances of VM occlusion.

### v4

@@ 4.1 Datasets and metrics
H3.6M [16]. We use (S1, S5, S6, S7, S8) for training and (S9, S11) for testing. As in [19, 7, 32, 33], we report MPJPE and PA-MPJPE for poses that are derived from the estimated meshes. We also report Mean Per Vertex Error (MPVE) for the whole mesh.
3DPW [54] is collected in natural scenes. Following the previous works [32, 33, 24, 61], we use the train set of 3DPW to learn the model and evaluate on the test set. The same evaluation metrics as H3.6M are used.
SURREAL [53] is a large-scale synthetic dataset with GT SMPL annotations and has diverse samples in terms of body shapes, backgrounds, etc. We use its training set to train a model and evaluate the test split following [7].
@@ 4.2 Implementation Details
We learn $64$ virtual markers on the H3.6M [16] training set. We use the same set of markers for all datasets instead of learning a separate set on each dataset. Following [19, 7, 39, 61, 26, 23, 33, 32], we conduct mix-training by using MPI-INF-3DHP [38], UP-3D [28], and COCO [34] training set for experiments on the H3.6M and 3DPW datasets. We adapt a 3D pose estimator [47] with HRNet-W48 [46] as the image feature backbone for estimating the 3D virtual markers. We set the number of voxels in each dimension to be $64$, i.e. $D=H=W=64$ for 3D heatmaps. Following [19, 26, 39], we crop every single human region from the input image and resize it to $256\times 256$. We use Adam [22] optimizer to train the whole framework for $40$ epochs with a batch size of $32$. The learning rates for the two branches are set to $5\times 10^{-4}$ and $1\times 10^{-3}$, respectively, which are decreased by half after the $30^{th}$ epoch. Please refer to the supplementary for more details.
@@
HMR [19] CVPR’18 - 85.1 73.6 55.4
BodyNet [52] ECCV’18 Skel. + Seg. 65.8 - -
GraphCMR [26] CVPR’19 3D vertices 103.2 87.4 63.2
SPIN [25] ICCV’19 - 82.3 66.7 43.7
DecoMR [62] CVPR’20 IUV image 68.9 52.0 43.0
Pose2Mesh [7] ECCV’20 3D skeleton 68.8 56.6 39.6
PC-HMR [37] AAAI’21 3D skeleton 59.8 51.7 37.9
DynaBOA [14] TPAMI’22 - 70.7 55.2 34.0
Ours Virtual marker 44.7 36.9 28.9
Table 3: Comparison to the state-of-the-arts on SURREAL [53] dataset. means training on the test split with 2D supervisions. “Skel. + Seg.” means using skeleton and segmentation together.
@@ 4.3 Comparison to the State-of-the-arts
Results on H3.6M. Table 2 compares our approach to the state-of-the-art methods on the H3.6M dataset. Our method achieves competitive or superior performance. In particular, it outperforms the methods that use skeletons (Pose2Mesh [7], DSD-SATN [49]), body markers (THUNDR) [61], or IUV image [62, 64] as proxy representations, demonstrating the effectiveness of the virtual marker representation.
@@
Results on 3DPW. We compare our method to the state-of-the-art methods on the 3DPW dataset in Table 2. Our approach achieves state-of-the-art results among all the methods, validating the advantages of the virtual marker representation over the skeleton representation used in Pose2Mesh [7], DSD-SATN [49], and other representations like IUV image used in PyMAF [64]. In particular, our approach outperforms I2L-MeshNet [39], METRO [32], and Mesh Graphormer [33] by a notable margin, which suggests that virtual markers are more suitable and effective representations than detecting all vertices directly as most of them are not discriminative enough to be accurately detected.
@@
Results on SURREAL. This dataset has more diverse samples in terms of body shapes. The results are shown in Table 3. Our approach outperforms the state-of-the-art methods by a notable margin, especially in terms of MPVE. Figure 1 shows some challenging cases without cherry-picking. The skeleton representation loses the body shape information so the method [7] can only recover mean shapes. In contrast, our approach generates much more accurate mesh estimation results.
@@
Table 4: Ablation study of the virtual marker representation for our approach on H3.6M and SURREAL datasets. “Skeleton” means the sparse landmark joint representation is used. “Rand virtual marker” means the virtual markers are randomly selected from all the vertices without learning. (c) is our method, where the learned virtual markers are used.
4.4 Ablation study
Virtual marker representation. We compare our method to two baselines in Table 4. First, in baseline (a), we replace the virtual markers of our method with the skeleton representation. The rest are kept the same as ours (c). Our method achieves a much lower MPVE than the baseline (a), demonstrating that the virtual markers help to estimate body shapes more accurately than the skeletons. In baseline (b), we randomly sample $64$ from the $6890$ mesh vertices as virtual markers. We repeat the experiment five times and report the average number. We can see that the result is worse than ours, which is because the randomly selected vertices may not be expressive to reconstruct the other vertices or can not be accurately detected from images as they lack distinguishable visual patterns. The results validate the effectiveness of our learning strategy.
@@ 4.4 Ablation study
Figure 4: Mesh estimation results of different methods on H3.6M test set. Our method with virtual marker representation gets better shape estimation results than Pose2Mesh which uses skeleton representation. Note the waistline of the body and the thickness of the arm.
@@
Figure 1 shows some qualitative results on the SURREAL test set. The meshes estimated by the baseline which uses skeleton representation, i.e. Pose2Mesh [7], have inaccurate body shapes. This is reasonable because the skeleton is oversimplified and has very limited capability to recover shapes. Instead, it implicitly learns a mean shape for the whole training dataset. In contrast, the mesh estimated by using virtual markers has much better quality due to its strong representation power and therefore can handle different body shapes elegantly. Figure 4 also shows some qualitative results on the H3.6M test set. For clarity, we draw the intermediate representation (blue balls) in it as well.
Figure 5: Visualization of the learned virtual markers of different numbers of $K=16,32,96$, from left to right, respectively.
@@
Number of virtual markers. We evaluate how the number of virtual markers affects estimation quality on H3.6M [16] dataset. Figure 5 visualizes the learned virtual markers, which are all located on the body surface and close to the extreme points of the mesh. This is expected as mentioned in Section 3.1. Table 5 (GT) shows the mesh reconstruction results when we have GT 3D positions of the virtual markers in objective (1). When we increase the number of virtual markers, both mesh reconstruction error (MPVE) and the regressed landmark joint error (MPJPE) steadily decrease. This is expected because using more virtual markers improves the representation power. However, using more virtual markers cannot guarantee smaller estimation errors when we need to estimate the virtual marker positions from images as in our method. This is because the additional virtual markers may have large estimation errors which affect the mesh estimation result. The results are shown in Table 5 (Det). Increasing the number of virtual markers $K$ steadily reduces the MPVE errors when $K$ is smaller than $96$. However, if we keep increasing $K$, the error begins to increase. This is mainly because some of the newly introduced virtual markers are difficult to detect from images and therefore bring errors to mesh estimation.
$K$ GT Det
@@
Table 5: Ablation study of the different number of virtual markers ($K$) on H3.6M [16] dataset. (GT) Mesh reconstruction results when GT 3D positions of the virtual markers are used in objective (1). (Det) Mesh estimation results obtained by our proposed framework when we use different numbers of virtual markers ($K$).
Figure 6: Mesh estimation comparison results when using (a) fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$, and (b) updated $\hat{\mathbf{A}}$. Please zoom in to better see the details.
Figure 7: Top: Meshes estimated by our approach on images from 3DPW test set. The rightmost case in the dashed box shows a typical failure. Bottom: Meshes estimated by our approach on Internet images with challenging cases (extreme shapes or in a long dress).
Coefficient matrix. We compare our method to a baseline which uses the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$. We show the quality comparison in Figure 6. We can see that the estimated mesh by a fixed coefficient matrix (a) has mostly correct pose and shape but there are also some artifacts on the mesh while using the updated coefficient matrix (b) can get better mesh estimation results. As shown in Table 6, using a fixed coefficient matrix gets larger MPVE and MPJPE errors than using the updated coefficient matrix. This is caused by the estimation errors of virtual markers when occlusion happens, which is inevitable since the virtual markers on the back will be self-occluded by the front body. As a result, inaccurate marker positions would bring large errors to the final mesh estimates if we directly use the fixed matrix.
@@
(a) Ours (fixed) 64.7 51.6
(b) Ours 58.0 47.3
Table 6: Ablation study of the coefficient matrix for our approach on H3.6M dataset. “fixed” means using the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ to reconstruct the mesh.